The area of a rectangle is 200 sq cm and one dimension is 25 cm. Find the length of a diagonal of the rectangle (to the nearest whole number).
Answer by checkley71(8403)

25x=200
x=200/25
x=8 for the other dimention of the rectangle.
Now you have a right triangle with sides of 25 & 8.
So we find the hypotenuse (diagonal) using the pythagorean equation.
a^2+b^2=c^2
8^2+25^2=h^2
64+625=h^2
689=h^2
h=sqrt689
h=26 cm is the length of the hypotenuse (diagonal).